What is the potential difference between two points in a circuit if 200J of electrical energy are changed to other forms of energy when 25 coulombs of electric charge pass? If the charge flows in 10 seconds, what is the current?
1
Expert's answer
2018-02-26T09:33:07-0500
We know, W = q.V …………………………(1)
Where , W =energy, q = charge and V = potential energy
Here, W = 200 J , q = 25 coulomb and we have to find V.
Now, put the value of W and q we get, 200 = 25 × V ………………………..(2)
Rearranging equation (2) we get,
V = 200/25 = 8 volts.
Again we know, I = q/t ………………….(3)
Where, q = charge, t = time, I = current
Here, q = 25 coulombs, t = 10 sec., we have to find I.
Now, put the value of q and t in equation (3), we get,
I = 25/10 = 2.5 amps.

Answer: Potential difference is 8 volts. and the current is 2.5 amps.